Solve: 2siny^2dx + xycos^2dy=0 ,
Given: y(2)=√π/2
1
Expert's answer
2020-09-13T18:11:12-0400

As per the question ,

Given differential equation is

"2sin(y^2)dx+xycos(y^2)dy=0"


Using seperation of variable method,

xy"cos(y^2)" "dy=-2sin(y^2)dx"

"\frac{ycos(y^2)}{sin(y^2)}dy=-\frac{2}{x}dx"

Integrating both the side "\int ycot(y^2)dy=-\int \frac{2}{x}dx" ....(1)

let y"^2=t"

Diffrentiate with respect to x

2ydy=dt

ydy="\frac{dt}{2}"

Substitute y "^2" =t in equation 1

"\int \frac{Cot tdt}{2}=-2logx+logc"

"\frac{log|sin t|}{2}=logx^{-2}+logc"

"log|sin (y^2)|^{0.5}=logx^{-2}+logc"

"log|sin (y^2)|^{0.5}=logx^{-2}c" { using loga+logb=logab}

"sin(y^2)^{0.5}=\frac{c}{x^2}"

"\sqrt{sin(y^2)}=\frac{c}{x^2}" ....(2)

at Y(2)="\sqrt{\frac{\Pi}{2}}"


"\sqrt{sin(\frac{\Pi}{2})}=\frac{c}{2^2}"

1="\frac{c}{4}"

c=4

Putting the value of c in equation 2

"\sqrt{sin(y^2)}=\frac{4}{x^2}"

This is the required solution.